Timetable Wednesday, September 11

Time  Room LG 1, HS 1  Room LG 1, HS 2  Room LG 1, HS 304

08.30 – 09.15 Plenary Talk (Chair: Kazimierz Malanowski) Irena Lasiecka: Control and optimization of hybrid PDE systems

09.20 – 09.50 (Chair: Marcus Wagner) Andrzej Cegielski: Residual selection in projection methods for convex minimization problems (Chair: Peter Fusek) Elena Resmerita: On total convexity, Bregman projections and stability in Banach spaces (Chair: Armin Hoffmann) Ivan Ginchev: On a second order condition in vector optimization

09.50 – 10.20 Robert Dylewski: Numerical behavior of projection methods with residual selection for convex minimization problems Aris Daniilidis: Subdifferential representation of convex functions: refinements and applications Valentin Gorokhovik: High order minimality conditions for an abstract vector optimization problem

10.20 – 10.45  Coffee break

10.45 – 11.30 Plenary talk (Chair: Marcin Studniarski) Dariusz Zagrodny: Per claim optimal reinsurance strategies

12.00  Excursion